Tyler has $60,000 to invest for 3 years. Estimate which investment option will earn more interest. calculate to verify your estimate.
A. 5.1% simple interest, calculated monthly
B. 4.9% compound interest, calculated annually

Answer by stanbon(75887)

Tyler has $60,000 to invest for 3 years. Estimate which investment option will earn more interest. calculate to verify your estimate.
A. 5.1% simple interest, calculated monthly
I = P*r*t
I = 60,000*0.051*3 = $9180
I = P(1+(r/n)^(nt) - P = 60000(1+0.051/12)^(12*3)-60000
=60000[(1+0.051/12)^(12*3)-1]
= 60000*0.164947 = $9896.83
--------------------------
Cheers,
Stan H.